Can someone help me on this?
I need to translate this statement to an algebraic equation. Let x represent the number in this problem:
The sum of two consecutive integers is 145. Find the two integers.
Thanks,
Ashley

Answer by checkley71(8403)

X+(X+1)=145 OR 2X+1=145 OR 2X=144 OR X=72 & X+1=73 OR 72+1=73 OR 73=73